OpenStudy (anonymous):

solve for the approximate value of x in the following expression lnx=3

OpenStudy (anonymous):

You'll have to take e of both sides :\[\ln e x= e^3 \] and ln e , cancels out so you'll get : \[x = e^3 \approx20.09\] Correct me if I'm wrong ^_^

OpenStudy (anonymous):

i still not get it